# Coding for Gaussian Two-Way Channels: Linear and Learning-Based Approaches

### v1

@@ VI-I Block-Length Gain of Two-Way Coding
TABLE III: Block-length gain of RNN-based (two-way) coding along different numbers of processing bits $K=K_{1}=K_{2}$ when $\text{SNR}^{\text{ch}}_{1}=1$dB and $\text{SNR}^{\text{ch}}_{2}=20$dB.
@@
Lastly, we investigate the block-length gain of two-way coding for GTWCs, where the block-length gain is measured as the sum-error, either sum-BLER or sum-BER. For two-way coding, we consider RNN-based coding with $r_{1}=r_{2}=1/3$. We consider two different channel scenarios: (i) $\text{SNR}^{\text{ch}}_{1}=1$dB and $\text{SNR}^{\text{ch}}_{2}=20$dB, and (ii) $\text{SNR}^{\text{ch}}_{1}=-1$dB and $\text{SNR}^{\text{ch}}_{2}=20$dB. First, in Table III, we investigate the sum-BER and sum-BLER performance under $\text{SNR}^{\text{ch}}_{1}=1$dB and $\text{SNR}^{\text{ch}}_{2}=20$dB by varying the numbers of processing bits $K=K_{1}=K_{2}$. We consider the sigmoid function for RNN-based coding. In the table, we introduce the measure “Sum-BLER ($L=60$)” to compare the BLER performances with different values of $K$, which is calculated as $(1-\text{BLER}_{1,K})^{L/K}+(1-\text{BLER}_{2,K})^{L/K}$, where $\text{BLER}_{i,K}$ is the BLER of block-length $K$ of User $i$. Both sum-BER and sum-BLER peak at $K=3$, implying that our RNN-based coding exploits the block-length gain up to $K=3$. We next investigate the different channel scenario, $\text{SNR}_{1}=-1$dB and $\text{SNR}_{2}=20$dB, in Table IV, where the softmax function is considered for RNN-based coding. In the table, sum-BER peaks at $K=4$, while sum-BLER peaks at $K=6$.
TABLE IV: Block-length gain of RNN-based (two-way) coding along different numbers of processing bits $K=K_{1}=K_{2}$ when $\text{SNR}^{\text{ch}}_{1}=-1$dB and $\text{SNR}^{\text{ch}}_{2}=20$dB.
@@
We observe that in the low SNR scenario, i.e., $\text{SNR}_{1}=-1$dB in Table IV, more numbers of bits for coding yields better sum-error performance than the case with the higher SNR, i.e., $\text{SNR}_{1}=1$dB in Table III. This is because two-way coding potentially benefits from noise-averaging effect by utilizing more numbers of channel uses under low SNR scenarios. Note that “Sum-BLER ($L=120$)” in Table IV represents the sum of BLERs with $L=120$, calculated similarly in Table III with $L=60$. For simulations of RNN-based coding, we have considered the two representative coding architectures based on the results of Tables  III and IV: $K=6$ bits with the softmax function and $K=3$ bits with the sigmoid function. It is important to note that, for specific SNR scenarios, the best sum-error performance is attained through (i) the appropriate selection of $K$ and (ii) a proper choice between sigmoid and softmax. As a result, given a certain channel environment, we can select the best-performing coding architecture from a set of pre-trained architectures that have been trained with various $K$ and activation functions.

### v2

@@ VI-I Block-Length Gain of Two-Way Coding
TABLE III: Block-length gain of linear (two-way) coding along different numbers of processing bits $K=K_{1}=K_{2}$ when $\text{SNR}^{\text{ch}}_{1}=1$dB, $\text{SNR}^{\text{ch}}_{2}=20$dB, and $r_{1}=r_{2}=1/3$.
$K=6$ 1.47E-1 5.07E-2 1.98E-1 6.72E-1 3.02E-1 1.97
$K=5$ 9.80E-2 1.20E-2 1.10E-1 4.60E-1 6.00E-2 1.52
$K=4$ 1.66E-2 3.66E-3 2.03E-2 6.66E-2 1.47E-2 8.44E-1
$K=3$ 5.51E-4 8.36E-4 1.39E-3 3.31E-3 5.08E-3 1.61E-1
$K=2$ 6.44E-5 2.70E-7 6.47E-5 1.30E-4 4.66E-7 3.90E-3
$K=1$ 7.76E-4 $<$1E-12 7.76E-4 7.76E-4 $<$1E-12 4.55E-2
TABLE IV: Block-length gain of RNN-based (two-way) coding along different numbers of processing bits $K=K_{1}=K_{2}$ when $\text{SNR}^{\text{ch}}_{1}=1$dB, $\text{SNR}^{\text{ch}}_{2}=20$dB, and $r_{1}=r_{2}=1/3$.
@@
Next, we investigate the block-length gain of two-way coding for GTWCs, where the block-length gain is measured as the sum-error, either sum-BLER or sum-BER. For two-way coding, we consider linear coding and RNN-based coding with $r_{1}=r_{2}=1/3$. We consider two different channel scenarios: $(\text{SNR}_{1},\text{SNR}_{2})=(1,20)$dB and $(\text{SNR}_{1},\text{SNR}_{2})=(-1,20)$dB. We have four results summarized in respective tables: (i) linear coding under $(\text{SNR}_{1},\text{SNR}_{2})=(1,20)$dB in Table III, (ii) RNN-based coding under $(\text{SNR}_{1},\text{SNR}_{2})=(1,20)$dB in Table IV, (iii) linear coding under $(\text{SNR}_{1},\text{SNR}_{2})=(-1,20)$dB in Table V, and (iv) RNN-based coding under $(\text{SNR}_{1},\text{SNR}_{2})=(-1,20)$dB in Table VI. Overall, we investigate the sum-BER and sum-BLER performance by varying the numbers of processing bits $K=K_{1}=K_{2}$. We introduce the measure “Sum-BLER ($L=60$)” or “Sum-BLER ($L=120$)” to fairly compare the BLER performances with different values of $K$, which is calculated as $1-(1-\text{BLER}_{1,K})^{L/K}+1-(1-\text{BLER}_{2,K})^{L/K}$, where $\text{BLER}_{i,K}$ is the BLER of $K$ bits of User $i$. For RNN-based coding, we consider the sigmoid function in (ii) and the softmax function in (iv).
For linear coding, both sum-BER and sum-BLER peak at $K=2$ in both channel SNR scenarios, $(\text{SNR}_{1},\text{SNR}_{2})=(1,20)$dB in Table III and $(\text{SNR}_{1},\text{SNR}_{2})=(-1,20)$dB in Table V. This indicates that, under these channel SNRs, up to $K=2$, the feedback benefit from an increasing number of channel uses $N=K_{i}/r_{i}$ outweighs the performance loss from increasing modulation order $2^{K}$. However, beyond $K=2$, the performance loss by increasing modulation order outweighs the feedback benefit. It is important to note that the best choice of $K$ varies with different channel SNRs. For example, while the best performing $K$ is $K=2$ when $(\text{SNR}_{1},\text{SNR}_{2})=(1,20)$dB as shown in Table III, the best performing $K$ when $(\text{SNR}_{1},\text{SNR}_{2})=(1,30)$dB is $K=3$, as shown in Fig. 5(5(a)).
For RNN-based coding, in the case of $(\text{SNR}_{1},\text{SNR}_{2})=(1,20)$dB in Table IV, both sum-BER and sum-BLER peak at $K=3$, implying that our RNN-based coding exploits the block-length gain up to $K=3$. In the case of $(\text{SNR}_{1},\text{SNR}_{2})=(-1,20)$dB, in Table VI, sum-BER peaks at $K=4$, while sum-BLER peaks at $K=6$. For RNN-based coding, we observe that in the low SNR scenario, i.e., $\text{SNR}_{1}=-1$dB in Table VI, more numbers of bits for coding yields better sum-error performance than the case with the higher SNR, i.e., $\text{SNR}_{1}=1$dB in Table IV. This is because two-way coding potentially benefits from noise-averaging effect by utilizing more numbers of channel uses under low SNR scenarios.
TABLE V: Block-length gain of linear (two-way) coding along different numbers of processing bits $K=K_{1}=K_{2}$ when $\text{SNR}^{\text{ch}}_{1}=-1$dB, $\text{SNR}^{\text{ch}}_{2}=20$dB, and $r_{1}=r_{2}=1/3$.
$K=8$ 2.40E-1 1.41E-1 3.81E-1 9.17E-1 7.60E-1 2.00
$K=7$ 2.12E-1 9.38E-2 3.06E-1 8.49E-1 5.68E-1 N/A
$K=6$ 1.98E-1 4.56E-2 2.44E-1 7.76E-1 2.73E-1 2.00
$K=5$ 1.01E-1 1.44E-2 1.15E-1 5.04E-1 7.20E-2 1.83
$K=4$ 3.66E-2 8.51E-3 4.51E-2 1.46E-1 3.40E-2 1.64
$K=3$ 1.29E-2 8.46E-4 1.37E-2 3.86E-2 2.54E-3 8.90E-1
$K=2$ 7.59E-3 1.56E-4 7.75E-3 1.52E-2 3.17E-4 6.20E-1
$K=1$ 1.52E-2 2.21E-8 1.52E-2 1.52E-2 2.21E-8 8.41E-1
TABLE VI: Block-length gain of RNN-based (two-way) coding along different numbers of processing bits $K=K_{1}=K_{2}$ when $\text{SNR}^{\text{ch}}_{1}=-1$dB, $\text{SNR}^{\text{ch}}_{2}=20$dB, and $r_{1}=r_{2}=1/3$.
@@
It is important to note that, for RNN-based coding, at specific SNR scenarios, the best sum-error performance is attained through (i) the appropriate selection of $K$ and (ii) a proper choice between sigmoid and softmax. As a result, given a certain channel environment, we can select the best-performing coding architecture from a set of pre-trained architectures that have been trained with various $K$ and activation functions.
VI-J Experimental Validations for Linear Coding
Figure 14: The max of power values across various numbers of channel uses $N$ when $\eta_{1}=\eta_{2}=10$, $\text{SNR}^{\text{ch}}_{1}=0$dB, and $\text{SNR}^{\text{ch}}_{2}=10$dB. Assuming that the exhaustive search yields optimal solutions, the optimality loss of our approach (by solving (46) instead of (44)) is observed to be less than 3% up to $N=6$.
We first discuss the potential of optimality loss of our linear coding strategy in Sec. III-D when solving (46). This loss may arise due to (i) the use of the weighted sum method for multi-objective optimization, (ii) the choice of the simplified ${\bf F}_{2}\in\mathcal{F}_{2}$ in (47), and (iii) the use of the iterative optimization strategy in Algorithm 1. To demonstrate how much the performance of our coding strategy deviates from optimal performance, we adopt an exhaustive search method as our benchmark that solves (44). In the exhaustive search, we conduct an exhaustive search for ${\bf g}_{1}$, ${\bf F}_{1}$, ${\bf g}_{2}$, and ${\bf F}_{2}$, to minimize the max of powers in (44) while satisfying the SNR constraints of Users 1 and 2. Fig. 14 demonstrates the objective value in (44), i.e., max of powers, across various numbers of channel uses, $N$, when $\eta_{1}=\eta_{2}=10$, $\text{SNR}^{\text{ch}}_{1}=0$dB, and $\text{SNR}^{\text{ch}}_{2}=10$dB. Since the exhaustive search incurs huge computational complexity, we could obtain results only for small values of $N$, specifically up to $N=6$. For $N=1,2$, the two users do not use any cooperation. When $N\geq 3$, the two users cooperate, thereby reducing the max of powers. We find that the performance gap between our coding strategy and the exhaustive search is less than 3% at each $N$. This implies that the benefit of analytical tractability from solving (46) instead of (44) might outweigh the low optimality loss in performance.
Figure 15: Convergence of the weighted sum power value through the iterations in our iterative optimization for linear coding, with $N=6$, $\alpha=0.7$, $\eta_{1}=\eta_{2}=10$, $\text{SNR}^{\text{ch}}_{1}=0$dB, and $\text{SNR}^{\text{ch}}_{2}=10$dB. The objective value decreases and converges to a stable value across the iterations.
Next, we discuss the convergence of the iterative optimization for linear coding in Algorithm 1, discussed in Sec. III-E. While this iterative process does not guarantee convergence, which is challenging to prove theoretically, we empirically demonstrate convergence through numerical experiments shown in Fig. 15. For simulations, we consider $N=6$, $\alpha=0.7$, $\eta_{1}=\eta_{2}=10$, $\text{SNR}^{\text{ch}}_{1}=0$dB, and $\text{SNR}^{\text{ch}}_{2}=10$dB We conduct 30 Monte Carlo simulations and average the objective value (the weighted sum of power values in (46)) at each iteration. Fig. 15 demonstrates that the objective value decreases and converges to a stable value across the iterations of the first and second sub-problems, respectively, in Sec. III-E1 and III-E2.
Figure 16: The max-power value, ${\max}\big{\{}\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{1}\|^{2}\big{]},\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{2}\|^{2}\big{]}\big{\}}$, obtained by solving the weighted sum problem in (46) across various $\alpha\in(0,1)$. The max-power value is shown to be unimodal across different values of $\text{SNR}^{\text{ch}}_{2}$, which experimentally supports the validity of the unimodality assumption.
Finally, we discuss the unimodality assumption made in Sec. III-F1. We have conducted numerical simulations in Fig. 16 to experimentally demonstrate the validity of the unimodality assumption. Fig. 16 depicts the objective value, ${\max}\big{\{}\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{1}\|^{2}\big{]},\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{2}\|^{2}\big{]}\big{\}}$ in (44), that is obtained by solving the weighted sum problem in (46) across various $\alpha\in(0,1)$. For simulations, we consider $\eta_{1}=10$, $\eta_{2}=10$, $\text{SNR}^{\text{ch}}_{1}=1$dB, and $N=10$. In Fig. 16, we observe that the max-power value is unimodal, i.e., it decreases monotonically up to a certain point (the minimizer) and then increases monotonically, across different values of $\text{SNR}^{\text{ch}}_{2}$, which experimentally supports the validity of the unimodality assumption.